A sink can be filled from the faucet in 9 minutes. It takes only 5 minutes to empty the sink when the drain is open. If the sink is full and both the faucet and the drain are open, how long will it take to empty the sink?
Answer by lwsshak3(11628)

A sink can be filled from the faucet in 9 minutes. It takes only 5 minutes to empty the sink when the drain is open. If the sink is full and both the faucet and the drain are open, how long will it take to empty the sink?
***
1/9=work rate of faucet
1/5=work rate of drain
let x=time it takes to empty the sink with both faucet and drain open
x/5-x/9=100% of job=1
9x-5x=45
4x=45
x=45/4=11-1/4 minutes